A chemist wants to make a 10% solution of fertilizer. How much water and how much of a 30% solution should the chemist mix to get 30L of a 10% solution?
Answer by oberobic(2304)

The key consideration with mixture problems is how much 'pure' stuff you need.
The chemist wants 30 liters of a 10% solution: 30 * .1 = 3 liters of pure fertilizer and 27 liters of water.
He has 30% solution on hand.
.
x = liters of 30% solution
.
30-x = liters of water to add, which is 0% pure fertilizer
.
.3*x + 0*(30-x) = .1*30
.
.3x = 3
.
x = 10 liters of 30% solution
.
30-x = 20 liters of pure water
.
Done.